OpenStudy (anonymous):

i need help with this: Suppose every edge of a cube measures 15 centimeters. What is the volume of the cube?

OpenStudy (anonymous):

V=a^3, where a is the length of the edge of cube.

OpenStudy (chaise):

The volume is equal to the base area times the height of the cylinder. The base area is (LxW) = 15x15. The height is equal to 15 also, cause it's square. 15x15x15=15^3 = 3375 = V Which is how the formula is derived. The total volume is 3375cm^3

OpenStudy (anonymous):

it exponents